# Predicate Invention from Pixels via Pretrained Vision-Language Models

### v1

@@ Abstract
Our aim is to learn to solve long-horizon decision-making problems in highly-variable, combinatorially-complex robotics domains given raw sensor input in the form of images. Previous work has shown that one way to achieve this aim is to learn a structured abstract transition model in the form of symbolic predicates and operators, and then plan within this model to solve novel tasks at test time. However, these learned models do not ground directly into pixels from just a handful of demonstrations. In this work, we propose to invent predicates that operate directly over input images by leveraging the capabilities of pretrained vision-language models (VLMs). Our key idea is that, given a set of demonstrations, a VLM can be used to propose a set of predicates that are potentially relevant for decision-making and then to determine the truth values of these predicates in both the given demonstrations and new image inputs. We build upon an existing framework for predicate invention, which generates feature-based predicates operating on object-centric states, to also generate visual predicates that operate on images. Experimentally, we show that our approach — pix2pred — is able to invent semantically meaningful predicates that enable generalization to novel, complex, and long-horizon tasks across two simulated robotic environments.
@@ 1 Introduction
A core research objective in robotics is to develop a robot that can make long-term decisions from low-level sensory inputs to accomplish a very broad range of tasks. Recent work has shown that model-free imitation learning can solve complex tasks directly from pixels (Chi et al. 2023; Zhao et al. 2023). However, sequential and compositional generalization to novel tasks beyond the training distribution remain open challenges, especially when task horizons are long and demonstration datasets are small. In this work, we take an aggressively model-based approach to address these challenges. In particular, we use sparse demonstration data to learn abstract world models (Silver et al. 2023; Liang et al. 2024) and plan in those world models to achieve strong compositional and sequential generalization.
Figure 1: Example train vs. test task in the “Burger” environment. Training tasks involve creating a single burger with a single ingredient or stacking two ingredients. Tasks at test time involve more objects, different object arrangements, and novel compositions of goals as compared to those at training time. For example, one of the tasks at test time is to create three burgers – two with multiple ingredients – and where the robot starts out holding something (no training task starts out with the robot holding something).
Figure 2: Pix2Pred overview. pix2pred are given an initial set of predicates, a set of controllers (temporally extended actions like pick, cook, etc.) and demonstrations of solving tasks in the form of goal-directed observation-action sequences. Our method learns more predicates and learns operators in terms of the learned and given predicates. Given a novel observation and goal at test time, we use a planner to generate a plan with our learned operators that achieves the goal. Goals in the train and test tasks are defined in terms of the given predicates; goals are written as text in the figure to simplify the figure.
As a simple, pedagogical example, consider the “Burger” domain depicted in Figure 1. The agent is given demonstrations showing how to cook a patty, how to cut lettuce, how to stack ingredients, and how to make a sandwich with a single ingredient. It is then tasked with making multiple burgers, each with a cooked patty and chopped lettuce, from novel initial images where some patties may be cooked already, some lettuce chopped already, and where the robot might start out holding an object. This test task is challenging for model-free imitation learning because it is much longer-horizon than the demonstrated tasks and involves novel numbers and configurations of objects. However, if the agent can learn to recognize a small number of abstract state features, such as when a particular patty is Cooked, when a particular head of lettuce is Chopped, and whether or not the robot is currently holding an object, – as well as abstract transition rules involving these concepts: cooking a patty causes it to go from Raw to Cooked, chopping lettuce causes it to go from Whole to Chopped, a robot can only pick up an object if it is not Holding something else – it can combine these concepts in many ways when planning and generalize substantially to new tasks.
We are specifically interested in learning PDDL models (Fox and Long 2003; McDermott et al. 1998) in the form of symbolic predicates (e.g., Cooked) and operators (e.g., cooking a patty causes it to go from Raw to Cooked) that can be leveraged by powerful classical planning algorithms (Helmert 2006). Such models feature several biases that are conducive to efficient planning and learning. Firstly, they are relational: world states and goals are characterized in terms of properties and relations between objects. Secondly, they are object-oriented: the number and properties of objects may vary across different problem instances, allowing the same model to support a wide range of specific tasks. Finally, they are sparse: each modeled action depends on and affects only a relatively small number of other objects, abstracting out irrelevant details and predicting core cause-effect interactions. Several recent works have demonstrated that such models can be learned efficiently to support generalization to tasks featuring new states, new goals, more objects, and longer time horizons than those demonstrated at training time (Silver et al. 2021, 2022, 2023; Asai and Fukunaga 2018; Ahmetoglu et al. 2022). However, missing from previous work is a method for learning predicates that can be computed from raw visual observations from very few demonstrations.
To address this gap, our first key idea is to use Vision-Language Models (VLMs) to propose and ground predicates. Given demonstrations, we leverage the commonsense knowledge of the VLM propose predicates with natural language names. Given an invented predicate and new states, we can again use the VLM to ground and evaluate the predicate directly on images. However, we find experimentally that the VLM’s proposed set of predicates do not support efficient planning or generalization across training tasks. Given this, our second key idea is to generate a large pool of candidates and subselect predicates by explicitly optimizing for efficient and effective planning. We extend a recent method for program-synthesis-based predicate invention (Silver et al. 2023) that searches over sets of predicates via hill-climbing to minimize a fast-planning objective. After running this procedure to subselect predicates, we obtain a compact set of predicates that supports efficient and effective downstream decision-making.
In experiments, we evaluate the extent to which our method — pix2pred — is able to learn symbolic predicates and operators that generalize to novel problems. Across four tasks in two simulated environments, we compare to several related baselines and find that pix2pred consistently achieves the highest success rate on novel problem instances involving more objects, more complex goals, and longer horizons than those demonstrated during training
@@ 2 Related Work
Decision-Making with Foundation Models. Our work is strongly inspired by impressive recently-demonstrated capabilities of Large (Vision) Language Models (LLMs, VLMs) across a variety of challenging text and image tasks (OpenAI 2023; Team et al. 2023; Dubey et al. 2024). We build on a large body of work that leverages foundation models for decision-making in robotics tasks (Ichter et al. 2023; Singh et al. 2023; Liang et al. 2022; Huang et al. 2022; Curtis et al. 2024; Huang et al. 2023, 2024; Hu et al. 2023a; Duan et al. 2024; Yang et al. 2024) (see Hu et al. (2023b) for a recent survey). However, many of these approaches operate under a number of different restrictive assumptions. Some are confined to a particular task distribution (e.g., only pick-and-place) or seek to synthesize new relatively short-horizon skills (e.g., pouring a liquid into a cup) (Huang et al. 2023, 2024; Duan et al. 2024). Others assume skills are provided, and attempt to use foundation models to compose them to solve longer-horizon tasks (Ichter et al. 2023; Singh et al. 2023; Liang et al. 2022; Huang et al. 2022; Curtis et al. 2024; Hu et al. 2023a; Quartey et al. 2024; Kumar et al. 2024a). These approaches all use a foundation model directly for planning, which has been shown to perform significantly worse than dedicated planning algorithms that use an appropriate abstraction (Valmeekam et al. 2022; Kambhampati et al. 2024). By contrast, our approach leverages foundation models to construct abstractions that are fed into a planning system at inference time. We treat the work of Hu et al. (2023a) as representative of previous approaches that use foundation models to plan, and compare directly to variants of this approach in our experiments (Section 5).
Learning Abstractions for Planning. We build on a long line of work that learns abstractions for efficient planning (Bertsekas, Castanon et al. 1988; Jong and Stone 2005; Abel, Hershkowitz, and Littman 2016; Ugur and Piater 2015; Asai and Fukunaga 2018; Konidaris, Kaelbling, and Lozano-Pérez 2018; Silver et al. 2021; James, Rosman, and Konidaris 2022; Mao et al. 2022; Chitnis et al. 2022; Kumar et al. 2023; Silver et al. 2023; Han et al. 2024; Liu et al. 2024). Several of these works assume that predicates or operators are already given (Chitnis et al. 2021; Kumar et al. 2023; Mao et al. 2022; Siegel et al. 2024) or focus on the problem of learning predicates and other abstractions from online interaction (James, Rosman, and Konidaris 2022; Liang et al. 2024), or assume access to dense natural language descriptions with each demonstration (Liu et al. 2024; Han et al. 2024). By contrast, we aim to invent predicates and operators in an offline fashion from image-based demonstrations with no supervision on the kinds or number of predicates to learn. Like us, “skills to symbols” (Konidaris, Kaelbling, and Lozano-Pérez 2018) performs predicate invention given skills, but operates on point cloud data and uses a bisimulation objective to select predicates that is different from the fast-planning objective that we use. Our work builds directly on the method from  Silver et al. (2023), which learns predicates and operators that operate over a handcrafted, low-dimensional feature space. We extend this approach to learn predicates and operators that operate directly on images, and we demonstrate experimentally (Section 5) that this is necessary to solve tasks where attributes of objects important for decision-making cannot be easily extracted and encoded into a low-dimensional feature space.
3 Problem Setting and Background
We consider the problem of learning from demonstrations in deterministic, fully-observed environments. pix2pred takes as input a set of core object-types with meaningful natural-language names, an initial set of predicates, a set of parameterized controllers, and demonstrations of solving tasks in the form of goal-directed observation-action sequences, where goals are expressed in terms of the initial predicates and the sequences are segmented by controller. Each observation contains raw sensor data that includes one or more images characterizing the world state at that time, as well as data from sensors on the robot. We assume access to a perception function that can identify basic attributes of objects (e.g., position, dimensions, etc.) from the raw sensor data and can annotate objects of the given object-types on the images with unique names. Each controller execution in the demonstration is labeled with its name, the names of the objects it operates on in terms of names given by the perception function, and the continuous parameters it was executed with. From this, pix2pred learns predicates and operators (specifically in the form of a PDDL domain file (McDermott et al. 1998)), as well as samplers that produce the continuous parameters for the controllers. Below, we describe the precise structure of our demonstrations more formally. We also provide background on symbolic predicates, operators, samplers, and an existing framework for predicate invention that we build on.
3.1 Demonstrations
Our demonstrations are provided within a set of training tasks drawn from some distribution $\mathcal{T}$. We are ultimately interested in learning to solve any task $T$ drawn from this distribution. All tasks occur within a common environment.
Environments and Tasks
We model an environment as a tuple $\langle I,A,f,\Lambda,P\rangle$, where $I$ is the observation space, $A$ is the action space, and $f$ is an unknown transition function, denoted $f:I\times A\rightarrow I$. An observation $\text{obs}\in I$ consists of a sequence of one or more images ($\text{obs}^{\text{img}}$) representing camera images from cameras operated by the robot, as well as a vector of continuous real-valued numbers ($\text{obs}^{\text{feat.}}$) representing raw sensor data associated with the robot, such as GPS and encoder data. The agent receives some observation $\text{obs}_{t}$ at some timestep $t$, and can execute an action $a\in A$. The environment will then return a new observation $\text{obs}_{t+1}\leftarrow f(\text{obs}_{t},a)$. We assume $f$ is deterministic. $\Lambda$ is a finite set of object types, where each type $\lambda\in\Lambda$ has a name (e.g., robot, patty, etc.) and a tuple of real-valued features representing basic attributes of the object (e.g., position for most objects, position and joint angles for the robot).
Within an environment, a task is a tuple $\langle\mathcal{O},i_{0},g\rangle$. $\mathcal{O}$ is a set of objects, each with a type from $\Lambda$. The environment’s perception function $P:I\times\mathcal{O}\rightarrow\mathcal{X}$ processes each observation into a structured state used for downstream decision-making. These states have two parts: an image-based state, ($x^{\text{img}}$), and a feature-based state, ($x^{\text{feat}}$), concatenated together to form a state $x\in\mathcal{X}$. The perception function gives a unique name to each object in the observations and annotates the images with their names, as shown in Figure 2; this forms the image-based state. In practice this can be achieved by combining open-vocabulary object-detection and segmentation models (Ren et al. (2024)). The perception function also identifies the basic attributes of each object according to its type; this forms the feature-based state. In practice, this can be implemented with custom detectors for each property of each object, and is simple when the attributes are basic properties like pose and dimensions. We assume that objects are fully-observable to the perception function (i.e., there is no observation in which certain objects or types of objects are unobservable). $i_{0}$ is the initial observation of the task, and $x_{0}$ represents the initial state of a task produced by the perception function, where $x_{o}\leftarrow P(i_{0},\mathcal{O})$. We will henceforth refer to a task as involving this initial state: $\langle\mathcal{O},x_{0},g\rangle$. We describe the goal $g$ in a the next section.
Initial Predicate Set and Goals
In order to specify the goal $g$ of a task, as well as provide a basis for decision-making and learning, we provide an initial set of predicates $\Psi_{\text{init}}$ as part of the demonstrations. Each predicate $\psi$ can be thought of as a template for a binary classifier over states. Specifically, a predicate is characterized by a name, an ordered list of types $(\lambda_{1},\dots,\lambda_{m})$ and a classifier function $c_{\psi}:\mathcal{X}\times\mathcal{O}^{m}\to\{\text{true},\text{false}\}$ parameterized by typed object arguments. A ground atom $\underline{\psi}$ consists of a predicate $\psi$ and objects $(o_{1},\dots,o_{m})$. Each ground atom induces a particular binary state classifier $c_{\underline{\psi}}:\mathcal{X}\to\{\text{true},\text{false}\}$, where $c_{\underline{\psi}}(x)\triangleq c_{\psi}(x,(o_{1},\dots,o_{m}))$. For example, the ground atom Cooked(patty0) would have the value “true” if the object patty0 is cooked. [^1] It would have this value if its classifier was accurate. We discuss what happens if this classifier is inaccurate in Section 4.3.
We assume the initial predicates $\Psi_{\text{init}}$ are sufficient for representing task goals, and may additionally contain concepts useful across all tasks in the environment, but are insufficient on their own in that additional predicates will be necessary to capture all concepts critical for solving tasks. Specifically, the goal $g$ of a task is a set of ground atoms involving predicates in $\Psi_{\text{init}}$ and objects in $\mathcal{O}$. A goal $g$ is said to hold in a state $x$ if for all ground atoms $\underline{\psi}\in g$, the classifier $c_{\underline{\psi}}(x)$ returns true. We use the term “goal predicates” for predicates that are used in defining the goal. We provide a discussion about the necessity of predicate invention despite having these predicates in Appendix A.4.
Actions as Hybrid Controllers

### v4

# From Pixels to Predicates: Learning Symbolic World Models via Pretrained Vision-Language Models

@@ Abstract
Our aim is to learn to solve long-horizon decision-making problems in complex robotics domains given low-level skills and a handful of short-horizon demonstrations containing sequences of images. To this end, we focus on learning abstract symbolic world models that facilitate zero-shot generalization to novel goals via planning. A critical component of such models is the set of symbolic predicates that define properties of and relationships between objects. In this work, we leverage pretrained vision-language models (VLMs) to propose a large set of visual predicates potentially relevant for decision-making, and to evaluate those predicates directly from camera images. At training time, we pass the proposed predicates and demonstrations into an optimization-based model-learning algorithm to obtain an abstract symbolic world model that is defined in terms of a compact subset of the proposed predicates. At test time, given a novel goal in a novel setting, we use the VLM to construct a symbolic description of the current world state, and then use a search-based planning algorithm to find a sequence of low-level skills that achieves the goal. We demonstrate empirically across experiments in both simulation and the real world that our method can generalize aggressively, applying its learned world model to solve problems with a wide variety of object types, arrangements, numbers of objects, and visual backgrounds, as well as novel goals and much longer horizons than those seen at training time.
Keywords: Planning, Learning from Demonstration, Vision-Language Models
@@ 1 Introduction
A core research objective in robotics is to make long-term decisions from low-level sensory inputs to accomplish a very broad range of tasks [36]. For instance, a useful household robot should be able to satisfy a variety of requests, such as “wipe and clear my dining table”, “make me a cup of healthy juice”, or even “cook me a protein-rich meal”, across a wide range of novel household environments. Even when provided with a library of chainable short-horizon parameterized skills (such as moveto, pick, place, wipe, turn-knob), solving such tasks remains challenging. The robot must not only select and sequence the right skills, but also ground each one with the correct parameters—identifying which objects to act upon, where to move, how to grasp, where to place, or how far to turn a knob. Critically, the space of possible variations in objects, configurations, visual backgrounds, and goal specifications is extremely large: even seemingly simple instructions can require substantially different skill sequences and parameters in different situations.
In this work, we choose to tackle these challenges in a learning-from-demonstration setting with an aggressively model-based approach. Given a handful of demonstrations that use known parameterized skills, we learn symbolic world models [42, 30] composed of abstract, interpretable properties and object relations (i.e., predicates) that ground directly into low-level sensory inputs (i.e., pixels). Because these models capture task-agnostic world dynamics, they can generalize across objects, embodiments (e.g., from human to robot), and tasks (e.g., from simple settings at training time to more complex, novel settings at test time). Additionally, the structure of the models affords efficient search-based task planning [33, 10] to solve novel tasks at inference time [14].
Figure 1: Overview of pix2pred in the Cleanup domain: Given 66 human demonstrations showcasing the effects of distinct skills (e.g., wiping, dumping) and a small initial predicate set, pix2pred invents new predicates (e.g., NoObjectsOnTop(?table)) and learns symbolic operators. At test time, it uses a search-based planner that operates over the learned model to solve a novel multi-step task in a visually distinct environment—e.g., retrieving an eraser from a bin, clearing an obstacle, wiping the table, and returning the eraser.
As an illustrative setting, consider the Cleanup domain shown in Figure 1. At training time, the robot receives a few egocentric visual demonstrations of a human sequencing available skills — such as grasping, dumping, and wiping — to accomplish two specific goals: wiping a table, and throwing away a particular object. At inference time, it must achieve a novel goal of wiping, and putting away the eraser afterward in an entirely different room where the eraser is initially in a bin (and thus the robot must first dump the bin before using the eraser and then putting it back in the bin). This sequencing is challenging for model-free imitation learning as it requires producing an entirely different action sequence in a scenario with novel objects, and backgrounds. However, if the agent can learn and reason over a small set of conceptual distinctions in the world state, including whether an object is an eraser or a surface is clear, then it can compose these via planning to solve unseen tasks.
Learning predicates for symbolic world models [25, 42, 13, 30] in this setting is challenging for two reasons. First, we are provided no initial information or explicit constraints on the predicates to be learned: we must decide the number of predicates, the types and numbers of objects they should operate over, as well as their implementation (i.e, how they process input visual information). Second, we must invent predicates that lead to a symbolic world model that generalizes well. There are an infinite number of possible world models that could describe the dynamics illustrated by a set of demonstrations: we wish to select one that enables efficient planning across the entire distribution of tasks the demonstration tasks are drawn from.
To address these challenges, we introduce a new predicate-invention method — pix2pred — that exploits the perceptual grounding and common-sense capabilities of Vision-Language Models (VLMs) within a program synthesis framework. Our method assumes as input demonstrations annotated by intuitively-named parameterized skills and objects and a small set of initial predicates primarily used to describe the demonstrations’ goals. From these, it generates a diverse pool of candidate predicates — both programmatic, feature-based predicates (as in [42]) and VLM-derived, pixel-grounded predicates that capture high-level, semantically meaningful concepts (e.g., NoObjectsOnTop(?table)). Given this pool, we run a subselection procedure that learns symbolic operators and selects the predicate and operator set that optimizes an efficient planning objective [42]. This generate-then-subselect procedure allows pix2pred to filter redundant or unreliable predicates and retain only those that accurately ground in pixels and are useful for downstream decision-making.
In experiments, we find that pix2pred is able to learn symbolic world models that operate over pixels from just a handful (<15<15) of egocentric human demonstrations (Figure 1). These models enable a real-world Boston Dynamics Spot robot to solve two complex, long-horizon tasks with different layouts, objects, and visual appearances. We leverage three simulated domains developed in different previous works to rigorously benchmark our approach’s performance against several ablations and baselines from the literature. From this, we find that pix2pred achieves the highest success rates on novel problem instances involving more objects, more complex goals, and longer horizons.
@@ 2 Related Work
Decision-Making with Foundation Models. Our work is inspired by recently-demonstrated capabilities of large (vision) language models (LLMs, VLMs) in a variety of challenging text and image tasks [35, 46]. We build on a large body of work that leverages foundation models for decision-making in robotics [20, 44, 29, 19, 7, 18, 17, 16, 9, 51, 39] (see Hu et al. [15] for a recent survey). However, these approaches operate under a number of different restrictive assumptions. Some are confined to a particular task distribution (e.g., only pick-and-place) or seek to synthesize new relatively short-horizon skills (e.g., pouring liquid into a cup) [18, 17, 9]. Others assume skills are provided, and attempt to use foundation models to compose them to solve longer-horizon tasks [20, 44, 29, 19, 7, 16, 37, 27]. We similarly assume skills are provided, but leverage foundation models to construct abstractions that are used by a search-based planning system at inference time. We treat the work of Hu et al. [16] as representative of previous foundation model planning approaches and compare directly to variants in our experiments (Section 5).
Figure 2: Labeling and proposal (Left) A small subset of the proposed predicates for the Cleanup domain. (Right) Truth values for these predicates (green is true and red is false) in a particular state, as determined by the VLM. This data is ultimately used to subselect the correct predicates.
Learning Abstractions for Planning. We build on a long line of work that learns abstractions for efficient planning [3, 22, 1, 47, 2, 25, 43, 41, 21, 34, 6, 26, 42, 13, 31, 38]. Several of these methods assume that predicates or operators are already given [5, 34, 32, 40] or focus on the problem of learning predicates and other abstractions from online interaction [21, 30], or assume access to dense natural language descriptions with each demonstration [31, 13]. By contrast, we aim to invent predicates and operators in an offline fashion from image-based demonstrations with no supervision on the kinds or number of predicates to learn.
Our approach takes strong inspiration from “skills to symbols” [25] and follow-up work [21], which also aim to learn symbolic world models from a set of skills. James et al. [21] in particular learns abstractions by estimating option initiation sets and effects from interaction data, partitioning options into subgoal types, and merging objects into effect-equivalence classes to form predicates. In contrast, our method differs in two key ways. First, we leverage a vision–language model to propose and ground predicates directly from raw demonstrations, rather than inducing them through state partitioning. This enables learning from just a few short, embodiment-agnostic demonstrations, while biasing the representation toward human-interpretable, “commonsense” predicates (e.g., IsEraser(), IsColorGreen). Second, instead of relying on effect-equivalence criteria for learning predicates, we explicitly optimize predicate selection with a planning-efficiency objective, ensuring that the learned world models are directly tailored for effective long-horizon planning.
Our work builds directly on the bilevel planning framework of Silver et al. [42], which learns predicates, operators, and samplers compatible with Task and Motion Planning (TAMP) [11]. However, unlike Silver et al., which assumes a handcrafted, noiseless object-centric state space, our approach introduces algorithmic modifications (e.g., soft precondition intersection) to handle the noise inherent in VLM-generated visual predicates. Closely related concurrent works, VisualPredicator [30] and SkillWrapper [52], also leverage VLMs for predicate invention but are unable to learn from offline embodiment-agnostic demonstrations. In particular, these approaches typically require the agent to execute exploratory skills and learn from contrastive signals between successful and failed skill or plan executions. Our method targets the learning-from-demonstration setting: it learns entirely from a small set of offline, embodiment-agnostic demonstrations (e.g., human videos) that contain only positive skill executions. This allows us to acquire models without requiring online interaction or environment resets, enabling us to validate our approach on real-world tasks with a Boston Dynamics Spot robot (Section 5) trained using human video demonstrations. Ultimately, we view our method as complementary to these online approaches [30, 52]: pix2pred could be used to initialize a world model from offline data, which is subsequently refined into a provably sound and complete model online by these methods.
3 Background and Problem Setting
In this section, we describe how we model states, actions, and objects, and then provide background on predicates and symbolic world models. We also describe the inputs and outputs of the training (learning) and inference (deployment) phases of our method.
3.1 States, Actions, and Objects
We wish to perform decision making given an image-based state $s_{t}$ at timestep $t$ given a set of user-specified objects $\mathcal{O}$. Each object $o\in\mathcal{O}$ consists of a name, an optional type (denoted $\lambda$), and a descriptor string. If a type is specified, it is associated with a continuous feature vector composed of object properties (e.g. pose, size). Following previous work [42, 28], we assume that the agent is equipped with feature detectors (e.g. pose estimation within a pre-specified map) corresponding to each of these properties. Descriptions are short user-provided natural-language phrases (such as “blue mop” or “small round coffee table”) that, in the context of a particular set of images, suffice to designate a specific object. Importantly, we assume throughout that these descriptors enable clear disambiguation and identification of the particular object within the image-based state, and the corresponding objects are present in the initial state.
The state $s_{t}$ itself is comprised of a set of images: $s^{\text{img}}_{t}=\{i^{0}_{t},i^{1}_{t},\ldots,i^{j}_{t}\}$ (potentially captured from multiple cameras) and an “object-centric” state $s^{\text{obj}}_{t}$, which includes the set of continuous feature vectors (e.g. pose, size) for every unique object with corresponding type that is in the images. Overall, $s_{t}\leftarrow s^{\text{img}}_{t}\mathbin{\|}s^{\text{obj}}_{t}$, where $\mathbin{\|}$ denotes concatenation. Thus, the overall state space $S$ is composed of an image-based state space, and a feature-based state space.